how do we do 3 (x + 5) = 2x + 20
Found 2 solutions by Alan3354, lynnlo:
Answer by Alan3354(69443)

how do we do 3 (x + 5) = 2x + 20
-------------
1st, say what you want to do.

Answer by lynnlo(4176)

3(x+5)=2x+20
3(5+x)=20+2x
15+3x=20+2x
subtract 2x from both sides
(3x-2x+15=(2x-2x+20
add like terms in 3x-2x
3x-2x=x
x+15=(2x-2x)+20
look for sums of zero
2x-2x=0
x+15=20
isolate terms with x to the left
subtract 15 from both sides
x+(15-15)=20-15
look for two terms with sum of zero
15-15=0
x=20-15
evaluate 20-15
20-15=5
x=5